OpenStudy (anonymous):

Gardener A can mow an entire lawn in 1 hour less time than Gardener B. If they work together, they can mow the lawn in 20/9 hours. How many hours will it take Gardener A to mow the lawn alone?

OpenStudy (radar):

Gardner B takes 5 hr Gardner A takes 4 hr.

OpenStudy (radar):

Let x =gardener A time in hours then x-1= gardener B time in hours. when working together they accomplished the mowing in 20/9 hours using the product over sum method: \[x(x-1)\over 2x-1\]=20/9 cross multiplying: 9x(x-1)=40x-20 \[9x ^{2}-49x+20=0\] use quadratic formula getting x=5, x=4

OpenStudy (radar):

verify: product over sum 5(4) 20 ---- = ---- 5+4 9

OpenStudy (radar):

Could of also used the reciprocal of the sums of reciprocal.

OpenStudy (radar):

1 20 ----------------------------- = ---- 1 1 9 -------- + --------- x x-1 This too will work out to be the same quadratic equation as above and the same answer.

OpenStudy (radar):

Any questions?

OpenStudy (anonymous):

thanks radar!

OpenStudy (radar):

u r welcome